OpenStudy (anonymous):

3(0.1-0.4x)= 0.18x+0.06 find X

OpenStudy (anonymous):

0.3 - 1.2x = 0.18x + 0.06 0.3 - 0.06 = 0.18x + 1.2x 0.24 = 1.38x 0.24 / 1.38 = x x= 0.173

OpenStudy (anonymous):

my book says x=0.8

OpenStudy (anonymous):

please check the question that you have written here.. if it is correct then there is no problem in the solution..

OpenStudy (anonymous):

No, I wrote it write

OpenStudy (anonymous):

then the solution is right too.. perfectly right..

OpenStudy (anonymous):

ok thanks :)